OpenStudy (anonymous):

if someone can help me i will give a medal and it will be the last question i ask

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

What is the product of the quantity m squared plus 2 times m minus 8 all over 6 times p • 12 times p all over 4 times m plus m squared? 2 times the quantity m minus 2 all over m, m ≠ −4, 0 and p ≠ 0 m times the quantity m minus 2 all over 72 times p squared, p ≠ 0 m times the quantity m minus 2 all over 72 times p squared, m ≠ −4, 0 and p ≠ 0 2 times the quantity m minus 2 all over m times p, m ≠ −4, 0 and p ≠ 0

OpenStudy (anonymous):

Setup the problem first: \[\frac{m^2+2m-8}{\frac {4m+m^2}{6p*12p}}\] When dividing fractions, any fractions, its easier to visualize if you flip the bottom fraction and multiply. Such as: \[{(m^2+2m-8) * \frac {(4m+m^2)}{6p*12p}} \]

OpenStudy (anonymous):

First fraction should be \[\frac{m^2+2m-8}{\frac{6p12p}{4m+m^2}}\]

OpenStudy (anonymous):

I made a mistake the first time

OpenStudy (anonymous):

So, then you multiply \[\frac{(m^2+2m-8)*(4m+m^2)}{72p^2}\] the top, and combine any like terms. Can you solve it from there?

OpenStudy (anonymous):

Yes. thank you. i was confused but thank u for breaking it down for me

OpenStudy (anonymous):

Yeah the wording is a bit confusing. Its easier to see with number :)